4. A police car chases a holdup car speeding at 180 km/h and 20 min ahead of time. Find the speed of the police car when it is to catching up with holdup car in 50 min.?
1
Expert's answer
2018-06-26T09:27:08-0400
d=vt=180(50/60)=150 km.
But
d=VT.
V=d/T=150/((50-20)/60)=300 km/h.